# Hamiltonian Dynamics of Fröhlich Condensates in Classical Systems

### v1

@@ Abstract
Despite the above, the impact of a constant coupling on a Fröhlich system was still investigated. In this scenario, our results showed that the interaction energy decreases exponentially as Fröhlich condensation takes place. More precisely, the interaction energy in the lowest frequency mode becomes of the same order of magnitude as the harmonic energy, then rapidly diverges into negative values (not shown). This result is clearly not compatible with our Hamiltonian model whereby $H_{int}$ is always considered a small perturbation as compared to $H_{0}$. Again, using Eqs (5) and (6) helped fix this issue by keeping the interaction energy reasonably low even when strong condensation occurs (Fig. S4, see also note [20]).
@@
Discussion and perspectives—The present study deals with the foundations of the condensation phenomenon proposed by Fröhlich in the 1960s aimed at explaining coherent excitations in biomolecular systems [4]. Motivated by recent experimental findings [2, 3], we simulated the dynamics of a general classical Hamiltonian that incorporates all the key elements of Fröhlich’s models. Our results show that condensation, including strong condensates with reasonable energy values, can emerge under specific types of nonlinear coupling, supporting the physical plausibility of Fröhlich’s effect. However, small deviations from the original model, for example, via the addition of resonances involving high-frequency modes of the bath or the use of a nonlinear potential without bath mediation, result in suppressing the phenomenon. These factors should be carefully examined when investigating the condensation effect in real biomolecular structures.
@@
Given the large number of free parameters involved in the dynamics of Fröhlich’s systems, a full exploration of the parameter space is beyond the scope of this letter. Although off-resonance coupling [21] or higher-order coupling mechanisms could bring more realistic characteristics to the model [22], a more practical approach would be to tune coupling parameters directly from standard MD force fields. Possible strategies include normal mode analysis (NMA) to generate the set of eigenvectors needed to transition from real space to the space of protein normal modes [23, 24]. In this case, not only normal frequencies of the protein and the heat bath could be deduced but eigenvectors may also be used to evaluate the force field function in the space of modes. This would allow coupling coefficients in Eq. (3) to be deduced and used in our model to verify whether condensation can indeed arise. Finally, such an approach should give further insights into the reasons why previous MD studies have been unsuccessful in detecting Fröhlich’s effect and if missing ingredients are required.

### v2

@@ Abstract
Despite the above, the impact of a constant coupling on a Fröhlich system was still investigated. In this scenario, our results showed that the interaction energy decreases exponentially as Fröhlich condensation takes place. More precisely, the interaction energy in the lowest frequency mode becomes of the same order of magnitude as the harmonic energy, then rapidly diverges into negative values (not shown). This result is clearly not compatible with our Hamiltonian model whereby $H_{int}$ is always considered a small perturbation as compared to $H_{0}$. Again, using Eqs (5) helped fix this issue by keeping the interaction energy reasonably low even when sizable condensation occurs (Fig. S6 in [15]).
Thermostats—Aside from nonlinear terms and the choice of coupling coefficients, thermostats used to maintain the bath and the energy source at the target temperature may also have a non-negligible impact on the emergence of condensation. The influence of the Langevin friction coefficient of the bath ($\gamma$) and the source ($\gamma_{S}$) was investigated by varying each individually (Table S1) or both simultaneously (Table S2) from 0.03 to 1 ps-1, while keeping other parameters fixed. Results show that condensation is overall robust and its extent can be modulated by varying $\gamma$ and/or $\gamma_{S}$, highlighting the importance of dissipation and injection rates on Fröhlich effect.
Rescaling study—To further assess the robustness of the condensation mechanism and the impact of the parameters on the timescale of the condensation, coupling coefficients were all rescaled by the same multiplicative factor $\kappa$ over two orders of magnitude. Parameters used to produce Fig. 3 were taken as a reference ($\kappa=1.0$). The $\kappa$ factor was then varied from $0.01$ to $1$ and applied simultaneously to $\phi$, $\xi$, and $\lambda$, while the friction coefficients $\gamma$ and $\gamma_{S}$ are left unchanged. Fig. S7 in [15] shows the condensation ratio $\rho$ as a function of $\kappa$, along with the corresponding condensation time. The latter was obtained by fitting the average energy of the lowest-frequency mode as a function of time. Fitting was performed using the function $A-B\exp{(-t/\tau)}$ from the moment the source is activated ($\xi\neq 0$) when the energy is at the equipartition value of 300 K. The condensation time was then defined as the fitted coefficient $\tau$ multiplied by $\ln(2)$ corresponding to the time to reach half of the plateau value. From our analysis, we see that $\rho$ strongly increases as $\kappa$ decreases, while the condensation time follows a similar trend but rises much more sharply at small $\kappa$ values.
Figure 5: Blue: condensation ratio $\rho$ as a function of the scaling factor $\kappa$ multiplying the coupling coefficients $\phi$, $\xi$, and $\lambda$ and the friction coefficients $\gamma$ and $\gamma_{S}$ simultaneously. $\kappa=1.0$ corresponds to the parameters used in Fig. 3. Except for $\phi$, $\xi$, $\lambda$, $\gamma$, and $\gamma_{S}$, all other parameters are the same as in Fig. 3 for all $\kappa$ values. Green: condensation time (see main text for details).
The same rescaling study was repeated by applying the $\kappa$ factor to both the friction coefficients $\gamma$ and $\gamma_{S}$, and to the coupling constants. Results are displayed in Fig. 5. As in the previous case, a steep increase of the condensation time is observed as $\kappa$ becomes smaller. Although $\rho$ also increases, a clear concave dependence is found in this case, with $\rho$ stabilizing at very small $\kappa$. This highlights conditions in which the condensation time can grow to very large values while keeping the condensed energy relatively stable.
@@
Discussion and perspectives—The present study deals with the foundations of the condensation phenomenon proposed by Fröhlich in the 1960s aimed at explaining coherent excitations in biomolecular systems [5]. Motivated by recent experimental findings [2, 3, 4], we simulated the dynamics of a general classical Hamiltonian that incorporates all the key elements of Fröhlich’s models. Our results show that condensation, including strong condensates with reasonable energy values, can emerge under specific types of nonlinear coupling (Fröhlich resonances), supporting the physical plausibility of Fröhlich’s effect. Conversely, the addition of resonances involving high-frequency modes of the bath (Lifshits resonances) or the use of a nonlinear potential without bath mediation, result in suppressing the phenomenon. Interestingly, condensation still emerges without specific resonance conditions provided that nonlinear interactions are driven by low-frequency bath modes. This was implemented by introducing a low-pass filter on the bath frequencies, where the threshold is the frequency at which the number of verified Lifshits resonances exceeds that of Fröhlich’s. These factors should be carefully examined when investigating the condensation effect in real biomolecular structures.
@@
Given the large number of free parameters involved in the dynamics of Fröhlich’s systems, a full exploration of the parameter space is beyond the scope of this letter. Although higher-order coupling mechanisms could bring more realistic characteristics to the model [25], a more practical approach would be to tune the order-three coupling parameters so far considered directly from standard MD force fields targeting a specific model protein. Possible strategies include normal mode analysis (NMA) to generate the set of eigenvectors needed to transition from real space to the space of protein normal modes [26, 27]. In this case, not only normal frequencies of the protein and the heat bath could be deduced but eigenvectors may also be used to evaluate the force field function in the space of modes. This would allow coupling coefficients in Eq. (3) to be deduced and used in our model to verify whether condensation can indeed arise. Furthermore, we note that energy condensation remains sizable for a quite large interval of source and bath friction coefficients, including typical values used in MD simulations. Finally, while our frequencies are representative of low-frequency modes in proteins, coupling constants values were essentially chosen to observe a well-formed condensate within a reasonable timescale, typically on the order of several hundreds of ns. Reducing the magnitude of the coupling constants in our Hamiltonian was found to delay the formation of condensates. Although at this stage one cannot draw a direct link between the timescales explored by a typical MD force field and those explored in our model, our findings suggest that the length of force field-based simulations could be a possible limiting factor in observing a clear signature of well-formed Fröhlich’s condensates.